OpenStudy (anonymous):

Find f(x) if f(0) = 9 and the tangent line at (x, f(x)) has slope 5x

OpenStudy (anonymous):

\(5x\) is not a slope

OpenStudy (aakashsudhakar):

Ah, calculus! f'(x) = 5x f(x) = 2.5x^2 + C f(0) = 9 2.5(0)^2 + C = 9 C = 9 Therefore, the equation of f(x) can be represented by \[f(x) = \frac{ 5x^2 }{ 2 } + 9\]

OpenStudy (anonymous):

thank you!